A student working in the laboratory prepared the following reactants:

5 mL of 0.005M Cd2+(aq)

10 mL of 0.007M SCN-(aq)

10 mL of 0.8M HNO3(aq)

These reagents were mixed and allowed to stand for 10 minutes. The concentration of Cd(SCN)+ in the resulting equilibrium mixture is found to be 5 x 10−4M. Calculate the initial concentration of Cd2+(aq).

Calculate the equilibrium concentration of Cd2+(aq).Calculate the initial concentration of SCN-(aq).

Calculate the equilibrium concentration of SCN-(aq).

Calculate the Keq for the reaction mixture
.
Answer #1

. Initial conc. of Cd²+ = no. of millimoles Total volume in ml -0.oos xs tanol - 25 ohl 0.001 M & Initial cone. of SCN - = 0.Now, Cd2+ + sen = Cod (scn) + = Kea = [Cd (SCN) + ] [Cd2+] [SCN] Sxtot - 5*109 x 23 x10-3 → Keq = 435 - 435